A stationary proton is accelerated through a potential difference of 83.5 kV. What speed did the proton acquire?

1
Expert's answer
2022-03-09T10:33:28-0500

The work done by the field on proton is by definition:


"W = eV"

where "e = 1.60\times10^{-19}C" is the charge of proton, and "V = 86.5\times 10^3V" is the potential difference. According to the work-energy theorem, this work is equal to the kinetic energy of the proton:


"W = K= \dfrac{mv^2}{2}"

where "m = 1.67\times 10^{-27}kg" is the mass of proton, and "v" is its speed. Thus, obtain:


"\dfrac{mv^2}{2} = eV\\
v = \sqrt{\dfrac{2eV}{m}} = \sqrt{\dfrac{2\cdot 1.6\times 10^{-19}\cdot 86.5\times 10^3}{1.67\times 10^{-27}}} \approx 4.07\times 10^6m/s"

Answer. "4.07\times 10^6m/s"